# Tuning without Peeking: Provable Privacy and Generalization Bounds for LLM Post-Training

### v2

@@ G.1 Proof of Thm˜2
G.2 Theorem on Data Extraction and Proof of Corollary˜2
Corollary˜2 is a direct consequence of the following Theorem.
@@ Appendix H Experimental Details
Here we provide details for the settings of the experiments in Sec.˜5. The licenses for all datasets used in this work can be found in Tab.˜5.
@@
For example, in Carlini et al. [64], the authors propose to model the distribution of model’s “confidence” with a gaussian distribution; they train shadow models to estimate the parameters $(\mu_{in},\sigma_{in},\mu_{out},\sigma_{out})$ and perform a Likelihood-ratio test:
@@
Therefore, if a learning method does not impact the loss of the model over each sample, it would be difficult impossible to perform any MIA that relies on statistical tests based on the loss. We thus propose the two following metrics as proxies for the ability of membership inference attacks to succeed:
$\frac{1}{\text{\scriptsize num of samples}}\sum_{sample}|\text{\scriptsize per\_token\_}\text{NLL}_{\text{new model}}\{sample\}-\text{\scriptsize per\_token\_}\text{NLL}_{\text{inital model}}\{sample\}|$ (A.15)
$\frac{1}{\text{\scriptsize num of samples}}\sum_{sample}|\text{\scriptsize per\_token\_}\text{NLL}^{CoT,a}_{\text{new model}}\{sample\}-\text{\scriptsize per\_token\_}\text{NLL}^{CoT,a}_{\text{inital model}}\{sample\}|$ (A.16)

### v3

@@ G.1 Proof of Thm˜2
G.2 Theorem on Data Extraction and Proof of Cor.˜2
Cor.˜2 is a direct consequence of the following Theorem.
@@ Appendix H Experimental Details
Here we provide details for the settings of the experiments in Sec.˜5. The licenses for all datasets used in this work can be found in Tab.˜5. Experiments were implemented using Lingua [134] and vLLM [135].
@@
For example, in Carlini et al. [66], the authors propose to model the distribution of model’s “confidence” with a gaussian distribution; they train shadow models to estimate the parameters $(\mu_{in},\sigma_{in},\mu_{out},\sigma_{out})$ and perform a Likelihood-ratio test, with $p$ the density function of Gaussian distribution and $\phi$ the logit scaling as defined in the LiRA paper $\phi(t)=\log(\frac{t}{1-t})$:
@@
Therefore, if a learning method does not impact the loss of the model over each sample, it would be difficult or impossible to perform any MIA that relies on statistical tests based on the loss. We thus propose the two following metrics as proxies for the ability of membership inference attacks to succeed:
$\frac{1}{\text{\scriptsize num of samples}}\sum_{sample}|\text{\scriptsize per\_token\_}\text{NLL}_{\text{new model}}\{sample\}-\text{\scriptsize per\_token\_}\text{NLL}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}initial} model}}\{sample\}|$ (A.15)
$\frac{1}{\text{\scriptsize num of samples}}\sum_{sample}|\text{\scriptsize per\_token\_}\text{NLL}^{CoT,a}_{\text{new model}}\{sample\}-\text{\scriptsize per\_token\_}\text{NLL}^{CoT,a}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}initial} model}}\{sample\}|$ (A.16)
@@ Appendix I Additional Experiments
Experiment 3 for different tasks.
In Tab.˜9 we report the results of running BBoxER on tasks beyond mathematical reasoning. We update Llama-3.1-8B on the ARC-Easy training split and evaluate on the ARC-Easy test split, and repeat the same procedure for ARC-Challenge and Hellaswag. We report the mean and standard deviation across five random seeds of the relative improvement, or difference of accuracy, compared to the base model. We notice that on ARC Easy and Hellaswag we get an improvement, while it is not the case on ARC Challenge. This likely stems from the base model’s performance across tasks. On ARC Easy and Hellaswag, the base model scores about 80%, so “retrofitting” can provide meaningful improvement. In contrast, on ARC-Challenge the base is about 50% (random is at 25%), and BBoxER is designed to provide additional improvement as a lightweight add-on, rather than heavy duty training: if the base is weak, perturbing parameters won’t meaningfully improve the model even if we notice slight improvement on the training split.
Table 9: Experiment 3 (full update, normalization layers) different tasks. Relative accuracy improvement using BBoxER run with D-CMA, 5 seeds.
BBoxER budget b ARC Easy ARC Challenge Hellaswag
150 0.01 $\pm$ 0.10 -0.14 $\pm$ 0.12 0.16 $\pm$ 0.07
300 0.13 $\pm$ 0.06 -0.21 $\pm$ 0.17 0.25 $\pm$ 0.11
600 0.18 $\pm$ 0.16 -0.05 $\pm$ 0.25 0.47 $\pm$ 0.06
1200 0.85 $\pm$ 0.16 0.24 $\pm$ 0.21 0.89 $\pm$ 0.13
@@
Appendix K BBoxER Execution Example
In this section we provide a step by step execution example of BBoxER (Alg.˜1), using OneFifth (Alg.˜2, (1+1)-ES) as the BBO algorithm and Full Layer Update on the normalization layers as our choice of modifiedmodified.
Input: seed $\omega$, dataset $D$, BBO algorithm aa=OneFifth, budget $b$
  • step 1:
    initialize the internal state by seeding the rng with $\omega$, and fixing the step size $\sigma$
  • step 2:
    sample a $z\sim\sigma\mathcal{N}(0,I)$ via $a.ask()$
  • step 3:
    update model weights $m_{i}=modified(m_{0},z)$, where every RMSNorm layer weights $w_{i}^{(j)}$ are updated with $w_{i}^{(j)}=w_{0}^{(j)}\times\exp{(0.01*z_{j})}$ and $z=(z_{1},\dots,z_{\text{number RMSNorm layers}})$
  • step 4:
    $k_{i}$ = $a.numCases()$ = 2 (we only compare with the best so far, therefore 2 possible choices)
  • step 5:
    $choice_{i}=a.compare((m_{best},m_{i}),D)$ this comparison is either done by majority vote in A/B testing or by comparing the performance of $m_{best}$ and $m_{i}$ over $D$ (note that only the comparison result is needed and not the actual performance of each model)
  • step 6:
    $a.tell(choice_{i})$ and $I_{i}=(I_{i-1},choice_{i})$ update the internal state of algorithm a : if $m_{i}$ is better ($choice_{i}=1$) then $\sigma=2*\sigma$, otherwise ($choice_{i}=0$) $\sigma=2^{-\frac{1}{4}}*\sigma$
    repeat steps 2 to 6 for b iterations